# Unlocking Chemical Insights: Superior Molecular Representations from Intermediate Encoder Layers

### v1

@@ Abstract
Pretrained molecular encoders have become indispensable in computational chemistry for tasks such as property prediction and molecular generation. However, the standard practice of relying solely on final-layer embeddings for downstream tasks may discard valuable information. In this work, we challenge this convention by conducting a comprehensive layer-wise analysis of five diverse molecular encoders across 22 ADMET property prediction tasks. Our results demonstrate that embeddings from intermediate layers consistently outperform final-layer representations. Specifically, using fixed embeddings from the optimal intermediate layers improved downstream performance by an average of 5.4%, reaching gains up to 28.6%. Furthermore, finetuning up to these intermediate layers yielded even greater average improvements of 8.5%, with performance increases as high as 40.8%, achieving new state-of-the-art results on several benchmarks. Additionally, a strong positive correlation between fixed embedding performance and finetuning outcomes supports an efficient evaluate-then-finetune approach, enabling identification of optimal layers with reduced computational cost. These findings highlight the importance of exploring the full representational depth of molecular encoders to achieve substantial performance improvements and computational efficiency. The code is made publicly available at https://github.com/luispintoc/Unlocking-Chemical-Insights.
@@ 1 Introduction
Deep learning has reshaped molecular science, where pretrained molecular encoders are essential tools for applications from virtual screening in drug discovery to designing novel materials with desired properties goh2018smiles2vecinterpretablegeneralpurposedeep ; 10.1093/bioinformatics/bty593 ; Xu2023 ; sypetkowski2024on . These encoders, built on architectures such as Transformers vaswani2023attentionneed and Graph Neural Networks (GNNs) NIPS2017_5dd9db5e ; kipf2017semisupervisedclassificationgraphconvolutional , learn rich representations of molecular structures that can be applied to a wide range of downstream predictive tasks. The common practice of extracting molecular representations from the final encoder layer, while simple and widely adopted, rests on the implicit assumption that this layer consistently provides the most informative and task-relevant features.
However, this assumption is increasingly challenged by findings from Natural Language Processing (NLP) and Computer Vision, where intermediate layers often encode richer, more generalizable, or task-specific information, leading to significant performance gains liu-etal-2019-linguistic ; jin-etal-2025-exploring ; gurnee2024languagemodelsrepresentspace ; valeriani2023geometryhiddenrepresentationslarge . Despite the unique data modalities (e.g. molecular graphs, 3D conformers) and distinct pretraining objectives inherent to cheminformatics, systematic layer-wise investigations of molecular encoders have been notably scarce. This leaves a critical question unanswered: are we overlooking superior molecular representations by adhering to the final-layer convention in chemistry?
Figure 1: Percentage improvement in test metric of the best performing intermediate layer relative to the final-layer when evaluating fixed embeddings. A "–" symbol indicates that the final-layer was better.
To address this, we propose and implement a comprehensive two-stage evaluation study. First, we assess the quality of fixed embeddings extracted from every layer of various pretrained molecular encoders. Second, we evaluate the performance when the encoder is finetuned up to each respective layer for specific downstream tasks. This dual approach allows us to understand both the intrinsic quality of layer-wise representations and their potential as effective priors for task-specific adaptation.
@@
    We conduct an empirical study evaluating fixed embeddings from five diverse molecular encoders across 22 absorption, distribution, metabolism, excretion, and toxicity (ADMET) tasks. We demonstrate that in over 81% of model-task combinations, an intermediate layer outperforms the conventionally used final-layer, yielding an average increase of 5.4% in downstream performance.
@@
    We further show that these advantages persist when the encoders are finetuned up to the corresponding intermediate layers for the same set of tasks, again observing that in over 71% of cases, intermediate layers lead to notable improvements over finetuning the complete encoder, with an average gain of 8.5%.
@@
    We establish a strong positive Pearson correlation between the performance of fixed embeddings and finetuning performance on a per-layer basis. This suggests a practical evaluate-then-finetune strategy: fixed embedding performance can serve as a reliable guide for selecting the best layer to finetune, reducing both computational cost and trial-and-error.
@@ 2.3 Surrogate Evaluators for Embedding Quality
Evaluating the quality of learned representations without resorting to computationally expensive end-to-end finetuning is crucial for efficient model development and analysis. This has led to the use of various surrogate models, which are typically lightweight and quick to train, to probe the information encoded in fixed embeddings. These surrogates provide an estimate of the downstream utility of the embeddings across different tasks.
Several classes of models are commonly employed for this purpose. Linear probes, such as logistic regression or linear regression, are frequently used to assess the linear separability of concepts within embedding spaces Tripepi2008 . While simple, they offer a baseline understanding of what information is readily accessible. More complex, yet still relatively lightweight, tree-based ensemble models such as Random Forests or Gradient Boosted Decision Trees (e.g., XGBoost chen_xgboost_2016 , LightGBM ke_lightgbm_2017 , CatBoost dorogush2018catboostgradientboostingcategorical ) are often effective off-the-shelf predictors when applied to fixed embeddings, providing strong performance with some hyperparameter tuning priyadarsini2024improvingperformancepredictionelectrolyte ; soares2025representing . In some cases, small Multi-Layer Perceptrons (MLPs) are also trained as probes, offering a balance between simplicity and non-linear capacity Li2023 ; Masood2025 . More recently, models leveraging in-context learning principles have demonstrated significant promise. TabPFN hollmann2023tabpfntransformersolvessmall , for example, built upon a transformer architecture and pretrained on vast amounts of synthetic tabular data, can make accurate predictions without requiring task-specific gradient updates or hyperparameter tuning.
These surrogate evaluators enable efficient assessment of learned embeddings by pairing powerful pretrained encoders with lightweight predictors. From linear models to in-context transformer models, they offer a scalable alternative to finetuning, balancing speed and predictive reliability for downstream evaluation.
@@ 3 Methods
This section details the molecular encoder architectures investigated, the benchmark datasets and evaluation metrics employed, and the experimental protocols for evaluating layer-wise representations through both fixed embedding analysis and finetuning.
@@ 3.1 Molecular Encoder Models
  • Orb Family rhodes2025orbv3atomisticsimulationscale : We investigated two specific variants from the Orb family of machine learning force fields. They do not rely explicitly on equivariant message passing architectures but incorporate invariances through their feature design and data augmentation. The variants we used are:
@@ 3.2 Benchmark and Evaluation Protocol
We evaluated molecular encoder representations on the TDCommons (TDC) benchmark Huang2021tdc , a curated suite of 22 ADMET-related tasks spanning both regression and classification settings. Each task requires the prediction of a single target, a specific physicochemical or biological property, using molecular structures as input. ADMET properties critically influence the development and safety profiles of pharmaceutical compounds; thus, maximizing representational quality has tangible benefits for reducing experimental costs and enhancing predictive accuracy in drug discovery.
@@
To address the challenge of generalizing to novel chemical scaffolds, TDC includes scaffold-based data splits for training, validation, and testing, enabling meaningful evaluation of a model’s capacity to extrapolate to new chemical structures. Following TDC benchmark guidelines, performance for each task was quantified using its designated metric. For regression tasks, this metric was either Mean Absolute Error (MAE) or Spearman Rank Correlation Coefficient, while for classification tasks, it was either the Area Under the Receiver Operating Characteristic curve (AUROC) or the Area Under the Precision-Recall Curve (AUCPR). Thus, our comprehensive evaluation framework employs these four performance metrics. Further details on each task are provided in Appendix A.
3.3 Evaluating Fixed Layer-wise Embeddings
@@
To evaluate the intrinsic quality of representations from different layers of each molecular encoder, we utilized the hidden states from every layer as fixed feature embeddings for each molecule in the TDC task datasets. To derive a single vector representation per molecule from each layer, we followed the conventions established in the original publications of the respective models. For instance, MolFormer uses mean pooling over token embeddings, while Uni-Mol utilizes the representation of the special [CLS] token. For GNN-based models such as Pos-EGNN and the Orb variants, mean pooling over all node embeddings was employed to obtain the graph-level representation.
@@
For every TDC task, we retrieved the embeddings produced by each individual layer, trained TabPFN on the scaffold-split training portion of those embeddings and their labels, generated predictions for the matching test embeddings, and scored them with the metrics from Section 3.2.
@@ 3.4 Finetuning with Layer-wise Representations
We also performed a comprehensive finetuning study across all pretrained molecular encoders and TDC tasks. For each encoder and task, we extracted molecular representations from individual encoder layers (using the same pooling strategy as Section 3.3) and fed them into a task-specific prediction head. This prediction head consisted of two linear layers: the first projected to a hidden dimension, followed by a SiLU activation ELFWING20183 , 10% dropout, and a final linear layer projecting to the output dimension.
For each specific encoder layer being evaluated on each task, key hyperparameters, namely the prediction head’s hidden dimension, selected from {32, 256}, and the learning rate for the AdamW optimizer loshchilov2019decoupledweightdecayregularization , selected from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4}, were optimized via an independent hyperparameter search. Hyperparameter ranges were selected based on preliminary studies and align with configurations commonly used in the original publications of the respective models, effectively capturing performance variance without excessive computational demands.
During training, both the parameters of the encoder up to and including the selected layer and those of the prediction head were jointly finetuned; any encoder layers beyond the selected layer were excluded. Models were trained for 50 epochs with a fixed batch size of 64 and the learning rate was linearly warmed up over the first 5% steps and decayed over the remaining steps.
Validation performance was monitored at each epoch, and the model checkpoint yielding the best validation metric was used for test set evaluation. Although we focus on a single seed due to computational constraints, future work could further strengthen robustness through repeated runs.
@@ 4 Results and Discussion
In this section, we present the empirical findings from our layer-wise evaluation of five pretrained molecular encoders across 22 TDC ADMET tasks. We first show that intermediate layer representations consistently outperform the final layer when used as fixed feature extractors with TabPFN. We then show that these advantages largely persist when the encoders are finetuned up to the selected layer. Finally, we establish a correlation between these two evaluation paradigms, suggesting an efficient strategy for selecting optimal layers for finetuning.
@@ 4.1 Superior Performance of Intermediate Layer Embeddings
To assess the quality of representations from different depths of the pretrained molecular encoders, we first utilized their hidden states as fixed feature embeddings for downstream prediction with TabPFN. The performance of each layer was compared against the conventional approach of using only the final-layer’s output.
Our findings, summarized in Figure 1, reveal that intermediate layers often provide substantial improvements over the final-layer. The heatmap illustrates the percentage increase in the test metric when using the best-performing intermediate layer versus the final-layer for each model-task pair. Darker green signifies greater improvement from an intermediate layer, while lighter shades or white denote comparable performance or cases where the final-layer was optimal. Detailed layer-wise performance plots for each model and task are available in Appendix C.
The analysis highlights striking benefits: in 81% of model-task combinations, an intermediate layer performed better than the final-layer. Selecting the optimal intermediate layer boosted performance by an average of 5.4% over the final-layer. In some model-task combinations, these gains were particularly significant, reaching up to 28.6%.
The choice of encoder influences these gains. MolFormer and Uni-Mol benefit the most from intermediate layers, with average improvements of 7.9% and 6.7%, respectively. Their deeper architectures include 12 and 15 encoder blocks, which likely provide more intermediate representations to exploit. In contrast, the comparatively shallow Pos-EGNN shows an average improvement of only 3.7%.
Task-specific benefits were also evident. For instance, cyp2c9-substrate-carbonmangels saw substantial improvements from intermediate layers, with average gains of 19.8%. Conversely, tasks like clearance-hepatocyte-az and ames showed limited advantage, with intermediate layers outperforming the final-layer in only two of the five models examined. Variability across tasks may be attributed to differences in complexity, data distributions, or intrinsic representational demands specific to each ADMET endpoint. Furthermore, while prior work found that the most informative representations in language and vision models often emerge midway through the model skean2025layerlayeruncoveringhidden , our investigation in molecular encoders did not identify a specific depth as universally optimal; rather, the ideal layer depth varied considerably depending on the specific model and ADMET task.
Crucially, leveraging intermediate layers also yielded new state-of-the-art (SOTA) results. Several of these surpassed benchmarks previously reported in notwell2023admetpropertypredictioncombinations : Pos-EGNN’s Layer 0 achieved an AUROC of 0.994 on hia-hou (vs. 0.989 from the prior SOTA), MolFormer’s Layer 10 reached 0.940 AUROC on pgp-broccatelli (vs. 0.938), and Uni-Mol’s Layer 3 attained 0.641 Spearman correlation on clearance-microsome-az (vs. 0.630). Furthermore, Pos-EGNN’s Layer 3 set a distinct SOTA on dili, with an AUROC of 0.942, improving upon the 0.925 reported in Turon2022.12.13.520154 . These examples underscore the significant performance gains achievable by exploring these internal representations.

### v2

# AI4Mat-NeurIPS 2025

@@ Abstract
Pretrained molecular encoders have become indispensable in computational chemistry for tasks such as property prediction and molecular generation. However, the standard practice of relying solely on final-layer embeddings for downstream tasks may discard valuable information. In this work, we first analyze the information flow in five diverse molecular encoders and find that intermediate layers retain more general-purpose features, whereas the final-layer specializes and compresses information. We then perform an empirical layer-wise evaluation across 22 property prediction tasks. We find that using frozen embeddings from optimal intermediate layers improves downstream performance by an average of 5.4%, up to 28.6%, compared to the final-layer. Furthermore, finetuning encoders truncated at intermediate depths achieves even greater average improvements of 8.5%, with increases as high as 40.8%, obtaining new state-of-the-art results on several benchmarks. These findings highlight the importance of exploring the full representational depth of molecular encoders to achieve substantial performance improvements and computational efficiency. The code will be made publicly available.
@@ 1 Introduction
Deep learning has reshaped molecular science, where pretrained molecular encoders are essential tools for applications from virtual screening in drug discovery to designing novel materials with desired properties goh2018smiles2vecinterpretablegeneralpurposedeep ; sypetkowski2024on ; doi:10.1021/acs.chemrestox.5c00018 . These encoders, built on architectures such as Transformers vaswani2023attentionneed and Graph Neural Networks (GNNs) NIPS2017_5dd9db5e ; kipf2017semisupervisedclassificationgraphconvolutional , learn rich representations of molecular structures that can be applied to a wide range of downstream predictive tasks. The common practice of extracting molecular representations from the final encoder layer, while simple and widely adopted, rests on the implicit assumption that this layer consistently provides the most informative and task-relevant features.
However, this assumption is increasingly challenged by findings from Natural Language Processing (NLP) and Computer Vision, where intermediate layers often encode richer, more generalizable, or task-specific information, leading to significant performance gains liu-etal-2019-linguistic ; jin-etal-2025-exploring ; gurnee2024languagemodelsrepresentspace ; valeriani2023geometryhiddenrepresentationslarge . Despite the unique data modalities (e.g. molecular graphs, 3D conformers) and distinct pretraining objectives inherent to cheminformatics, systematic layer-wise studies of molecular encoders have been notably scarce. This leaves a critical question unanswered: are we overlooking superior molecular representations by adhering to the final-layer convention in chemistry?
(a) Tokenized-molecule Entropy
(b) Adjacent-layer CKA
Figure 1: Left: Tokenized-molecule entropy rises with depth but falls at the final block for MolFormer, Uni-Mol and PosEGNN, indicating compression; Orb models maintain higher spread. Right: Adjacent-layer CKA shows small changes across interior layers and a pronounced last-step change for most models. Depth is normalized from first encoder block (0%) to last (100%).
Our work consists of a two-stage analysis across five diverse pretrained encoders. First, we map the information flow through their hidden layers, and second, we empirically evaluate the performance of these layer-wise representations on a suite of downstream tasks.
@@
    We use two label-free probes, tokenized-molecule entropy and adjacent-layer linear centered kernel alignment (CKA), to characterize information flow within molecules and across model depth. Our analysis reveals that while the middle layers remain stable and capture general, transferable features, the pronounced change in the final layer signals both information compression and specialization for the pretraining task.
@@
    We conduct an empirical study evaluating frozen embeddings across 22 absorption, distribution, metabolism, excretion, and toxicity (ADMET) tasks. We demonstrate that in over 81% of model-task combinations, an intermediate layer outperforms the conventionally used final-layer, yielding an average increase of 5.4% in downstream performance.
@@
    We further show that these advantages persist when the encoders are partially finetuned. For the same set of tasks, we observe that finetuning up to an optimal intermediate layer leads to notable improvements over finetuning the complete encoder in over 71% of cases, with an average gain of 8.5%.
@@ 2.3 Surrogate Evaluators for Embedding Quality
Evaluating the quality of learned representations without resorting to computationally expensive end-to-end finetuning is crucial for efficient model development and analysis. This has led to the use of various surrogate models to probe the information encoded in frozen embeddings. These surrogates provide an estimate of the downstream utility across different tasks.
Linear probes, such as logistic regression or linear regression, are frequently used to assess the linear separability of concepts within embedding spaces Tripepi2008 . More complex, tree-based ensemble models such as Random Forests or Gradient Boosted Decision Trees (e.g., XGBoost chen_xgboost_2016 , LightGBM ke_lightgbm_2017 , CatBoost dorogush2018catboostgradientboostingcategorical ) are often effective when applied to frozen embeddings priyadarsini2024improvingperformancepredictionelectrolyte ; soares2025representing . In some cases, small Multi-Layer Perceptrons (MLPs) are also trained as probes Li2023 ; Masood2025 . More recently, models leveraging in-context learning principles have demonstrated significant promise. TabPFN hollmann2023tabpfntransformersolvessmall , built upon a transformer architecture and pretrained on vast amounts of synthetic tabular data, can make accurate predictions without requiring task-specific gradient updates or hyperparameter tuning. These surrogate evaluators enable efficient assessment of learned embeddings by pairing powerful pretrained encoders with lightweight predictors. From linear models to in-context transformer models, they offer a scalable alternative to finetuning, balancing speed and predictive reliability for downstream evaluation.
@@ 3 Methods
This section details the molecular encoder architectures investigated, the benchmark datasets and evaluation metrics employed, the information flow analysis and the experimental protocols for evaluating layer-wise representations through both frozen embeddings and finetuning.
@@ 3.1 Molecular Encoder Models
  • Orb Family rhodes2025orbv3atomisticsimulationscale : We investigated two specific variants from the Orb family of machine learning force fields. They do not rely explicitly on equivariant message passing architectures but incorporate invariances through their data augmentation. The variants we used are:
@@ 3.2 Benchmark and Evaluation Protocol
We evaluated molecular encoder representations on the TDCommons (TDC) benchmark Huang2021tdc , a curated suite of 22 ADMET-related tasks spanning both regression and classification settings. Each task requires the prediction of a single target, a specific physicochemical or biological property, using molecular structures as input. ADMET properties critically influence the development and safety profiles of pharmaceutical compounds; thus, maximizing representational quality has tangible benefits for reducing experimental costs and enhancing predictive accuracy in drug discovery doi:10.1021/acs.chemrestox.5c00273 ; MacDermottOpeskin2025ChemRxiv .
To address the challenge of generalizing to novel chemical scaffolds, TDC provides scaffold-based data splits that partition each dataset into training, validation, and testing sets. This ensures that compounds in the test set have distinct core structures from those in the training set, enabling a rigorous assessment of a model’s capacity to extrapolate to unseen chemical space.
@@
We aggregated the training partitions across tasks to construct an unlabeled corpus for the information flow analysis. For the empirical experiments, we used the prescribed splits per task without modification. Following TDC benchmark guidelines, performance for each task was quantified using its designated metric. For regression tasks, this metric was either Mean Absolute Error (MAE) or Spearman Rank Correlation Coefficient, while for classification tasks, it was either the Area Under the Receiver Operating Characteristic curve (AUROC) or the Area Under the Precision-Recall Curve (AUCPR). Thus, our comprehensive evaluation framework employs these four performance metrics. Further details on each task are provided in Appendix A.
3.3 Information Flow Probes
@@
We use two label-free probes to diagnose how information evolves across layers and to test the hypothesis that the final-layer is often over-specialized for pretraining. Let $x_{i}$ denote a molecule and $\ell$ an encoder layer. The token matrix is $H_{i,\ell}\in\mathbb{R}^{T_{i}\times d}$, where $T_{i}$ is the number of valid tokens/atoms for molecule $x_{i}$ and $d$ is the hidden dimensionality of layer $\ell$. To derive a single vector representation per molecule from each layer, we followed the conventions established in the original publications of the respective models. For instance, MolFormer uses mean pooling over token embeddings, while Uni-Mol utilizes the representation of the special [CLS] token. For GNN-based models such as Pos-EGNN and the Orb variants, mean pooling over all scalar node embeddings was employed to obtain the graph-level representation.
Tokenized-molecule entropy.
We first quantify the diversity of within-molecule token signals with a matrix-based entropy giraldo2014measuresentropydatausing . For each $x_{i}$ and layer $\ell$, we form the token Gram $K_{i,\ell}=H_{i,\ell}H_{i,\ell}^{\top}$, take its eigenvalues $\{\lambda_{t}\}$, normalize $p_{t}=\lambda_{t}/\sum_{u}\lambda_{u}$, and average the Shannon entropy of this spectrum across molecules:
$\mathrm{TME}(\ell)=\frac{1}{|\mathcal{D}|}\sum_{i\in\mathcal{D}}\Big(-\sum_{t}p_{t}\log p_{t}\Big).$
Higher values indicate that token embeddings are spread across many principal directions, expressing greater variety and lower redundancy. Lower values indicate that variance concentrates in a small number of directions, reflecting higher redundancy and representational compression, a collapse towards a low-rank subspace.
Adjacent-layer CKA on pooled molecule vectors.
We assess changes in the pooled space by comparing adjacent layers using linear centered kernel alignment (CKA) kornblith2019similarityneuralnetworkrepresentations . For two adjacent layers $\ell$ and $\ell{+}1$, we form $X$ by stacking, row-wise, the pooled molecule vectors from layer $\ell$ (one row per molecule), and likewise $Y$ from layer $\ell{+}1$. We then center each matrix across examples by subtracting the corresponding row-mean vector, yielding $\tilde{X}$ and $\tilde{Y}$. The centered second moments are then defined as
$S_{xx}^{(\ell)}=\tilde{X}^{\top}\tilde{X},\qquad S_{yy}^{(\ell+1)}=\tilde{Y}^{\top}\tilde{Y},\qquad S_{xy}^{(\ell,\ell+1)}=\tilde{X}^{\top}\tilde{Y}$.
And the linear CKA between layers is
$\mathrm{CKA}_{\ell\to\ell+1}=\frac{\|S_{xy}^{(\ell,\ell+1)}\|_{F}^{2}}{\|S_{xx}^{(\ell)}\|_{F}\,\|S_{yy}^{(\ell+1)}\|_{F}}\in[0,1],$
Values of $\mathrm{CKA}_{\ell\to\ell+1}$ close to 1 indicate that consecutive layers produce highly similar molecule embeddings, while values closer to 0 indicate larger geometric changes between layers. We report adjacent-layer CKA across depth to quantify, in a label-free and scale-free manner, how representation changes attenuate or persist throughout the encoder.
3.4 Evaluating Frozen Layer-wise Embeddings
To assess the predictive utility of representations from different layers, we use the hidden states from every encoder layer as frozen molecule embeddings for each TDC task. As explained in Section 3.3, we obtained a single vector representation for each molecule at every layer by following the conventions outlined in the original publications of the corresponding models.
@@ 3.4 Evaluating Frozen Layer-wise Embeddings
For every TDC task, we retrieved the embeddings produced by each individual layer, trained TabPFN on the scaffold‑split training portion of those embeddings and their labels, generated predictions for the matching test embeddings, and scored them with the metrics from Section 3.2.
@@ 3.5 Finetuning with Layer-wise Representations
We also performed a comprehensive finetuning study across all pretrained molecular encoders and TDC tasks. For each encoder and task, we extracted molecular representations from individual encoder layers, using the same pooling strategy as Section 3.4, and fed them into a task-specific prediction head. This prediction head consisted of two linear layers: the first projected to a hidden dimension, followed by a SiLU activation ELFWING20183 , 10% dropout, and a final linear layer projecting to the output dimension.
For each specific encoder layer being evaluated on each task, key hyperparameters, namely the prediction head’s hidden dimension selected from {32, 256} and the learning rate for the AdamW optimizer loshchilov2019decoupledweightdecayregularization selected from {1e-5, 2e-5, 5e-5, 1e-4, 2e-4}, were optimized via an independent hyperparameter search. Hyperparameter ranges were selected based on preliminary studies and align with configurations commonly used in the original publications of the respective models, effectively capturing performance variance without excessive computational demands.
During training, both the parameters of the encoder up to and including the selected layer and those of the prediction head were jointly finetuned; any encoder layers beyond the selected layer were excluded. Models were trained for 50 epochs with a fixed batch size of 64 and the learning rate was linearly warmed up over the first 5% steps and decayed over the remaining steps. Validation performance was monitored at each epoch, and the model checkpoint yielding the best validation metric was used for test set evaluation.
@@ 4 Results and Discussion
In this section, we begin with a label-free analysis of representation dynamics using the probes in Section 3.3. These depth-wise diagnostics reveal how token diversity and layer-to-layer geometry evolve inside each encoder. We then validate these findings by evaluating model performance across 22 TDC ADMET tasks in two settings. We first show that intermediate layer representations consistently outperform the final-layer when used as inputs to TabPFN. We then show that these advantages largely persist when the encoders are finetuned up to an optimal intermediate layer. Finally, we establish a correlation between these two evaluation paradigms, suggesting an efficient strategy for selecting optimal layers for finetuning.
4.1 Unsupervised Probes of Representation Dynamics
Figure 1 summarizes depth-wise behavior across models using two label-free probes. First, an analysis of token diversity using tokenized-molecule entropy reveals a pattern of end-of-encoder compression in several models. MolFormer and Uni-Mol exhibit a sharp drop in entropy in their final-layer, with Pos-EGNN showing a milder version of this effect. This drop indicates that the token representations collapse into a low-rank subspace, concentrating their variance into fewer principal directions. Such behavior is consistent with the final-layer becoming highly tailored to its pretraining objective. Conversely, both Orb variants maintain a broader spread across principal directions, indicating that token information stays more distributed instead of collapsing at the end.
Second, geometric stability between adjacent layers, captured by CKA, reveals that Uni-Mol sits near 1 across depth, indicating very similar pooled molecule embeddings from one block to the next. The remaining encoders live in the 0.90 to 0.95 range, which points to small but meaningful updates as depth increases. Most models then show a pronounced last-step drop in CKA signaling a large geometric remapping right before the output, that aligns with the pretraining task specialization and coincides with the token-level compression above. Viewed jointly, these trends support the view that, regardless of the specific architecture or pretraining objective, the final-layer is often over-specialized and does not always carry the most transferable features.
Figure 2: Percentage improvement in test metric of the best intermediate layer relative to the final-layer when evaluating frozen embeddings. Positive values mean the best non-final layer improves over the final-layer. Negative values mean the final-layer outperforms the best non-final layer.
@@ 4.2 Superior Performance of Intermediate Layer Embeddings
We evaluate downstream utility by using each layer’s hidden states as fixed feature embeddings with TabPFN. Figure 2 reports, for each model and task pair, the signed percent change of the best non-final layer relative to the final-layer, where green values indicate gains over the final-layer and red values indicate that the final-layer is better.
Across all evaluations, 81% of model and task combinations prefer a non-final layer. Negative values are present, although they are small in magnitude; when the final-layer wins, its average margin over the best non-final layer is about 1%. The choice of encoder influences these gains. MolFormer and Uni-Mol benefit the most from intermediate layers, with average improvements of 7.9% and 6.7%, respectively. In contrast, the comparatively shallow models show an average improvement of only $\sim$4%.
Moreover, while prior work in language and vision reports that the most informative representations often emerge midway through the model skean2025layerlayeruncoveringhidden , our empirical study did not identify a universally optimal depth for molecular encoders. Instead, the ideal layer varies by architecture and ADMET task, as illustrated by the full performance curves in Appendix C. This pattern aligns with the information flow findings. Across the interior of each encoder, adjacent-layer CKA is high and nearly constant, indicating that successive layers produce very similar pooled molecule vectors; in practice this means many intermediate layers supply comparably good features, so no single depth consistently dominates. In contrast, between the final two layers we observe a pronounced CKA drop, and tokenized-molecule entropy shows a concurrent decrease, both consistent with a last-stage geometric remapping and increased compression tailored to the pretraining objective. In aggregate, these observations explain why intermediate layers frequently outperform the final-layer while the precise winning depth varies by model–task pair.
Task-specific benefits were also evident. For instance, cyp2c9-substrate-carbonmangels saw substantial improvements from intermediate layers, with average gains of 19.8%. Conversely, tasks like clearance-hepatocyte-az and ames showed limited advantage, with intermediate layers outperforming the final-layer in only two of the five models examined. Furthermore, using intermediate layers also yielded new state-of-the-art (SOTA) results, for example: Pos-EGNN Layer 0 on hia-hou (AUROC 0.994 vs. 0.989), Uni-Mol Layer 3 on clearance-microsome-az (Spearman 0.641 vs. 0.630), and Pos-EGNN Layer 3 on dili (AUROC 0.942 vs. 0.925) notwell2023admetpropertypredictioncombinations ; Turon2022.12.13.520154 .